OpenStudy (e.mccormick):

@shamil98 Seen the matrix method for that?

OpenStudy (shamil98):

I haven't actually.

OpenStudy (e.mccormick):

L = S + M - 4 L + S + M = 70 2S = L-9 L - S - M = - 4 L + S + M = 70 -L + 2S+0M =-9 \(\left[\begin{array}{ccc|c} 1 & -1 & -1 & -4\\ 1 & 1 & 1 & 70\\ -1 & 2 & 0 & -9 \end{array}\right]\) Add R1 and R2 new R1 \(\left[\begin{array}{ccc|c} 2 & 0 & 0 & 66\\ 1 & 1 & 1 & 70\\ -1 & 2 & 0 & -9 \end{array}\right]\) R1 \(\times \dfrac{1}{2}\) \(\left[\begin{array}{ccc|c} 1 & 0 & 0 & 33\\ 1 & 1 & 1 & 70\\ -1 & 2 & 0 & -9 \end{array}\right]\) R2 + (-1)R1 new R2 R3 + R1 new R1 \(\left[\begin{array}{ccc|c} 1 & 0 & 0 & 33\\ 0 & 1 & 1 & 37\\ 0 & 2 & 0 & 24 \end{array}\right]\) \(R3 \(\times \dfrac{1}{2}\) \(\left[\begin{array}{ccc|c} 1 & 0 & 0 & 33\\ 0 & 1 & 1 & 37\\ 0 & 1 & 0 & 12 \end{array}\right]\) R2 + (-1)R3 \(\left[\begin{array}{ccc|c} 1 & 0 & 0 & 33\\ 0 & 0 & 1 & 25\\ 0 & 1 & 0 & 12 \end{array}\right]\) Longest = 33cm Medium = 25cm Shortest = 12cm

OpenStudy (shamil98):

What do R1 R2 R3 values mean?

OpenStudy (e.mccormick):

Row 1, 2, and 3.

OpenStudy (e.mccormick):

matrices use RC notation, Row by Column. So position 1,2 is row 1, column 2. The R and C then get reused a lot of places.... but some books like M and N.

OpenStudy (shamil98):

Why is R1 multiplied by 1/2? I never took algebra 2 lol.

OpenStudy (e.mccormick):

To get the pivot points to 1. The left hand side wants 1 in each column, known as the pivots. If you have onely one pivot at 1 in each row, then the right hand side is the solution to the system.

OpenStudy (e.mccormick):

Well, one in each row and one in each column, I should say.... so one value at 1 for each.

OpenStudy (shamil98):

Just a question where does the 66 come from ?..

OpenStudy (e.mccormick):

70+(-4) Here is a more detailed explanation, but only 4 pages so reasonably short: www.epcc.edu/tutorialservices/valleverde/Documents/Gauss-Jordan_Method.pdf It includes another worked example.

OpenStudy (e.mccormick):

Was gonna share this with that guy after you worked the solution with them,.... but that question sort of died. LOL. So thought I would at least share it with you so you could see another way.

OpenStudy (shamil98):

Yeah, he never replied haha

OpenStudy (e.mccormick):

Matrices can also be really helpful in solving some calculus problems. Another neat tool from Linear Algebra is the determinant, which can be used to easily find the area of an ellipse, volume of a parallelopiped, see if there are solutions to complex systems, and so on.